OpenStudy (anonymous):

If the length of the legs in a right triangle are 7 and 8, what is the approximate length of the hypotenuse?

OpenStudy (phi):

do you know the pythagorean theorem ? \[a^2 + b^2 = c^2\] where a and b are the legs and c is the hypotenuse

OpenStudy (anonymous):

Just set up an equation using the pythagorean theorem i like this version of it to find the hypotenuse:\[c = \sqrt{a^2+b^2}\] where c is the hypotenuse and a and b are the legs. so: \[c = \sqrt{7^2+8^2}\] Simplify:\[c=\sqrt{49+64}\] becomes:\[\sqrt{113}=c\] The hypotenuse is: \[\sqrt{113}\]